OpenStudy (aonz):

hi, Can someone just please explain the process of how to work this inequation/absolute value question please?

OpenStudy (aonz):

I know how to work this out |2x-5| + |x+1| =8 but not these |5-x| - |3+2x| >1

OpenStudy (aonz):

@jim_thompson5910 help please :)

jimthompson5910 (jim_thompson5910):

solve |5-x| - |3+2x| =1 and tell me what you get

OpenStudy (aonz):

here is my was of working out this question for the first question. |2x-5| + |x+1| =8

jimthompson5910 (jim_thompson5910):

no need for the if part after the solution, you can just say the solutions are x = 4 or x = -4/3

OpenStudy (aonz):

but is the way i did this pretty much exactly the same with inequalities?

jimthompson5910 (jim_thompson5910):

sort of...you basically solve the equation for x first that will give you a list of values which will help you set up a sign chart

jimthompson5910 (jim_thompson5910):

which you can then use to solve the inequality

OpenStudy (aonz):

for the 3 test points i got